OpenStudy (anonymous):

A tank has a supply pipe P and an exhaust pipe Q. The pipe P takes 4 minutes less time to fill the tank than the pipe Q to empty it. When both are open, the tank is filled in 24 minutes. Find the time needed for the pipe P to fill the tank when Q is closed.

OpenStudy (phi):

I would say the tank is emptied at a rate of \[ \frac{1 \text{ tank}}{M \text{ mins}} \] and is filled at a rate of \[ \frac{1 \text{ tank}}{(M-4) \text{ mins}} \] Now use rate * time = # of tanks filled (1 in this case) if we both fill and empty the tank at the same time the combined rate is \[ \frac{1}{M-4} - \frac{1}{M} \] if we do this for 24 minutes, we get a filled tank: \[ 24\left(\frac{1}{M-4} - \frac{1}{M}\right) = 1 \] can you solve for M ?

OpenStudy (anonymous):

@phi M = -8, 12 Since there is no negative time, 12 minutes is the answer...?

OpenStudy (phi):

We have to be careful. I picked M to be the time needed to empty the tank (Probably should have defined M to be the time to fill the tank). so the tank empties in 12 minutes. The tank fills at a rate of \[ \frac{1}{M-4} = \frac{1 \text{ tank}}{8 \text{ minutes} } \] that means it takes 8 minutes to fill the tank. or, using algebra \[ \frac{1 \text{ tank}}{8 \text{ minutes} }\cdot t = 1 \text{ tank} \\ t = 1 \cancel{\text{ tank}}\cdot \frac{8 \text{ minutes} }{1 \cancel{\text{ tank}}}\\ t= 8 \text{ minutes} \]

OpenStudy (anonymous):

@phi Thank you!

OpenStudy (phi):

yw